OpenStudy (anonymous):

find the difference of two cubes: 128x^3-250y^3

Directrix (directrix):

Here's the factoring pattern for the sum or difference of two cubes. Let's see if we can apply it to your problem.

Directrix (directrix):

128 and 250 are not "perfect cubes" so factor 2 from the given expression. 128x^3-250y^3 = 2( 64 x³ - 125 y³ ) = 2* [ (4x)³ - (5y)³ ] = @ktawneil See if you can write the next step. Okay?